# Quantum-Corrected Thermodynamics of Conformal Weyl Gravity Black Holes: GUP Effects and Phase Transitions

### v1

@@ VII Gravitational Redshift in CWGBH Geometry
The gravitational redshift parameter $z$, which quantifies the fractional change in photon frequency between emission and observation points, can be systematically linked to the underlying metric structure through the well-established relation [49]
$z=\left[1-2\dot{r}\right]^{1/2}-1.$ (51)
This formula encapsulates the fundamental physical principle that gravitational redshift arises from the difference in gravitational potential between emission and observation points, with the metric function serving as the mathematical representation of this potential difference in the curved spacetime geometry.
Substituting the explicit form of the CWGBH metric function into this expression and performing the necessary algebraic manipulations, we obtain the complete analytical expression for gravitational redshift in CWG geometry
$z=\sqrt{-1+6\beta\gamma+\frac{2\beta(2-3\beta\gamma)}{r}-2\gamma r-2kr^{2}}-1.$ (52)
This remarkable expression reveals the complex interplay between all CWG parameters in determining the redshift characteristics. The presence of terms involving $\beta\gamma$ coupling, linear $\gamma r$ contributions, and quadratic $kr^{2}$ dependencies demonstrates how the higher-order nature of CWG field equations manifests in observable electromagnetic phenomena. Unlike the simple $1/r$ dependence characteristic of Schwarzschild redshift, the CWG result exhibits a rich radial structure that could potentially be detected through high-precision spectroscopic observations of radiation emanating from the vicinity of massive compact objects.
Figure 8 presents a comprehensive visualization of the gravitational redshift behavior in CWGBH geometry, illustrating how the redshift parameter $z$ varies as a function of radial distance and the various CWG parameters. The plot reveals several distinctive features that distinguish CWG redshift from its GR counterpart: the presence of oscillatory behavior at intermediate distances due to the linear $\gamma r$ term, the modified asymptotic behavior at large distances governed by the cosmological parameter $k$, and the enhanced redshift effects near the horizon arising from the complex parameter couplings. These features provide potentially observable signatures that could serve as discriminating tests between CWG and standard GR, offering new avenues for testing alternative gravity theories through electromagnetic observations of compact astrophysical objects such as neutron stars, stellar-mass BHs, and supermassive BHs in galactic centers.
Figure 8: Gravitational redshift $z$ in CWGBH geometry showing the complex radial dependence arising from CWG corrections. The plot demonstrates how the redshift behavior deviates from standard Schwarzschild predictions due to the presence of linear $\gamma r$ and quadratic $kr^{2}$ terms, providing potentially observable signatures for distinguishing CWG from conventional GR through precision spectroscopic measurements.
VIII Conclusion
@@
In this comprehensive theoretical investigation, we systematically explored the thermodynamic and quantum features of BHs arising from the Mannheim-Kazanas solution within the framework of CWG. Unlike the conventional Einstein-Hilbert paradigm based on second-order field equations, CWG employs a fourth-order action with local conformal symmetry, leading to fundamentally modified gravitational dynamics and significantly richer vacuum structures [66, 94]. The inclusion of additional potential terms, specifically the linear $\gamma r$ and quadratic $kr^{2}$ contributions in the metric function given by Eq.(10), enabled an alternative theoretical description of gravitational phenomena spanning both galactic and cosmological scales without requiring the introduction of dark matter components.
@@
Our analysis commenced with the systematic application of the Hamilton-Jacobi tunneling formalism to compute the Hawking temperature of CWGBHs, as detailed in Section III. This investigation revealed explicit contributions from all conformal parameters $\beta$, $\gamma$, and $k$, culminating in the comprehensive temperature expression presented in Eq.(26). The resulting thermal spectrum deviates markedly from that of standard Schwarzschild-AdS BHs, particularly due to the influence of the linear potential term $\gamma r$, demonstrating the remarkable sensitivity of BH radiation to conformal corrections [95, 96]. Table 1 provided quantitative evidence of these effects, showing temperature enhancements exceeding 75,000% in certain parameter regimes compared to baseline Schwarzschild values.
We subsequently incorporated quantum gravitational effects through the GUP framework, deriving the QC Hawking temperature expression that accounts for minimal length scale effects. This analysis demonstrated that the presence of a fundamental length scale systematically suppresses the thermal spectrum, particularly in the near-Planckian regime where $r_{h}\to l_{p}$. The resulting temperature suppression reflects a gravitational redshift in the emitted radiation and suggests the potential emergence of remnant-like structures during the final stages of BH evaporation [97, 98]. This quantum modification provides crucial insights into the transition between semiclassical and fully quantum gravitational regimes.
@@
In Section V, we employed an exponentially corrected entropy model, as expressed in Eq.(33), to systematically compute the complete set of QC thermodynamic potentials for CWGBHs. These calculations yielded detailed expressions for internal energy (Eq.(35)), Helmholtz free energy (Eq.(37)), pressure (Eq.(39)), enthalpy (Eq.(41)), and heat capacity (Eq.(44)). Figure 1 illustrates the complex energy landscape, revealing how scale-dependent corrections become increasingly important in the small horizon radius regime. Most significantly, Figure 6 demonstrated the existence of thermodynamic phase transitions through the heat capacity analysis, showing distinct regions of negative and positive heat capacity that correspond to thermodynamically unstable and stable phases, respectively.
@@
The JTE analysis presented in Section VI provided a sophisticated probe of isenthalpic processes in CWGBH thermodynamics within the extended phase space formalism. The behavior of the Joule-Thomson coefficient $\mu_{J}$ revealed distinct cooling and heating phases that depend sensitively on both the horizon radius and the CWG parameters. Figure 7 captured the QC phase structure, showing how the scale-dependent parameter $\gamma$ controls the location of Joule-Thomson inversion points and introduces increasingly complex thermal response behavior. These results demonstrated that CWG naturally accommodates rich phase structures that are absent in conventional GR treatments.
@@
Our investigation of gravitational redshift within CWGBH geometry, detailed in Section VII, revealed that the frequency shift of electromagnetic radiation emitted near these objects encodes valuable information about the underlying conformal metric structure. The derived redshift expression showed complex radial dependence arising from the interplay of all CWG parameters, providing potentially observable signatures that could serve as discriminating tests between CWG and standard GR through precision spectroscopic observations. Figure 8 shows how the redshift behavior deviates systematically from Schwarzschild predictions, offering new avenues for experimental tests of alternative gravity theories.
Throughout our analysis, we demonstrated that CWG offers a remarkably consistent and theoretically rich framework for studying BH thermodynamics beyond the limitations of GR, particularly when augmented with quantum corrections that become essential near the Planck scale. The intricate interplay between higher-derivative gravitational dynamics and quantum-statistical effects opened promising new avenues toward understanding quantum gravity phenomenology in astrophysically relevant contexts [99]. Our results provide both theoretical foundations and practical computational tools for investigating quantum gravitational effects in realistic BH systems.
@@
Looking toward future research directions, our comprehensive theoretical framework suggests several promising avenues for extended investigation. First, the development of more sophisticated numerical techniques for solving the complete nonlinear CWG field equations would enable precise predictions of observational signatures, including gravitational wave emission patterns, accretion disk spectra, and electromagnetic radiation characteristics that could be compared with data from current and future astronomical facilities. Second, the extension of our thermodynamic analysis to rotating CWGBHs would provide crucial insights into the effects of angular momentum on QC thermal properties, potentially revealing new classes of phase transitions and instabilities relevant to realistic astrophysical scenarios. Third, comprehensive observational studies involving gravitational lensing phenomena, particularly strong lensing effects around supermassive BHs, could provide direct constraints on the CWG parameters $\beta$, $\gamma$, and $k$ through comparison with high-resolution imaging data from facilities such as the Event Horizon Telescope. Fourth, the investigation of gravitational wave signatures from CWGBHs during merger events could yield distinctive waveform characteristics that differ from GR predictions, offering new opportunities for testing alternative gravity theories through advanced LIGO/Virgo observations and future space-based detectors.
Finally, another study of cosmological implications of CWG, particularly its role in explaining dark matter phenomena and cosmic acceleration without exotic matter components, might represent an extension of our BH-focused investigation. Therefore, our future research agenda is set to enhance our grasp of quantum gravity phenomena and offer empirical methods for examining the essential characteristics of spacetime geometry within extreme astrophysical environments.
Acknowledgements.
We express gratitude to Dr. Ercan Kılıçarslan and Dr. Heeseung Zoe for their constructive feedback. İ. S. appreciates the academic support from EMU, TÜBİTAK, ANKOS, and SCOAP3, as well as the networking opportunities through COST Actions CA22113, CA21106, and CA23130.

### v2

@@ 7 Gravitational Redshift in CWGBH Geometry
An important observational signature of CWGBHs is the gravitational redshift experienced by photons emitted near the event horizon. For static spherically symmetric spacetimes with line element $ds^{2}=-B(r)dt^{2}+B(r)^{-1}dr^{2}+r^{2}d\Omega^{2}$, the gravitational redshift between an emitter at radius $r_{e}$ and an observer at radius $r_{o}$ is given by
$z=\sqrt{\frac{B(r_{o})}{B(r_{e})}}-1.$ (57)
It is crucial to note that the MK metric is not asymptotically flat when $\gamma\neq 0$ or $k\neq 0$, since the metric function $B(r)$ diverges as $r\to\infty$ due to the linear $\gamma r$ and quadratic $kr^{2}$ terms. Consequently, the observer must be placed at a finite radius $r_{o}$ rather than at spatial infinity. This formula correctly reduces to $z=0$ in flat space ($\beta=\gamma=k=0$, where $B(r)=1$), ensuring physical consistency and yielding purely real values for all parameter combinations satisfying $B(r_{e})>0$ and $B(r_{o})>0$.
$\beta$ $\gamma$ $k$ $r_{h}$ $r_{e}=3r_{h}$ $z$
0.5 0 0 1.000 3.000 0.2186
1.0 0 0 2.000 6.000 0.2124
0.5 0.01 0 0.998 2.993 0.7002
1.0 0.01 0 1.990 5.970 0.6695
0.5 0 0.0001 1.000 3.000 0.7266
1.0 0 0.0001 1.999 5.998 0.7189
0.5 0.01 0.0001 0.997 2.992 1.0854
1.0 0.01 0.0001 1.989 5.968 1.0484
Table 4: Gravitational redshift $z$ for CWGBHs with observer at $r_{o}=100$ and emitter at $r_{e}=3r_{h}$ (photosphere region). The horizon radius $r_{h}$ is determined self-consistently from $B(r_{h})=0$. The first two rows correspond to the Schwarzschild limit ($\gamma=0$, $k=0$), while subsequent rows demonstrate the significant enhancement induced by the CWG parameters.
Table 4 presents the gravitational redshift for various CWG parameter combinations, with the emitter located in the photosphere region at $r_{e}=3r_{h}$ and the observer at $r_{o}=100$. In the Schwarzschild limit ($\gamma=0$, $k=0$), the redshift takes values $z\approx 0.21$$0.22$, consistent with standard GR predictions. The introduction of CWG parameters dramatically enhances the gravitational redshift. For $\beta=1$ and $\gamma=0.01$, the redshift increases from $z=0.2124$ to $z=0.6695$, representing a $215\%$ enhancement; this substantial increase arises because the linear potential term $\gamma r$ in $B(r)$ grows with radius, amplifying the ratio $B(r_{o})/B(r_{e})$. The cosmological parameter produces comparable effects: for $\beta=1$ and $k=0.0001$, the redshift rises to $z=0.7189$, a $238\%$ increase over the Schwarzschild value. When both $\gamma$ and $k$ are nonzero, their effects combine constructively; for $\beta=1$, $\gamma=0.01$, and $k=0.0001$, the redshift reaches $z=1.0484$, exceeding unity and representing a $394\%$ enhancement compared to Schwarzschild.
To further quantify the CWG corrections, we compare the redshift values for fixed emission and observation radii ($r_{e}=6$, $r_{o}=100$, corresponding to $\beta=1$ in the Schwarzschild limit) while varying $\gamma$:
$\gamma$ $k$ $z$ $\Delta z$
0 0 0.2124
0.001 0 0.2677 $26.0\%$
0.005 0 0.4634 $118.1\%$
0.010 0 0.6672 $214.1\%$
0.020 0 0.9911 $366.6\%$
Table 5: Comparison of gravitational redshift for Schwarzschild ($\gamma=0$) versus CWG BHs with varying $\gamma$. Fixed parameters: $\beta=1$, $k=0$, $r_{e}=6$, $r_{o}=100$. The percentage deviation $\Delta z$ is calculated relative to the Schwarzschild value.
As shown in Table 5, even modest values of the scale-dependent parameter $\gamma$ produce substantial redshift enhancements. For $\gamma=0.02$, the redshift approaches unity ($z\approx 0.99$), nearly five times the Schwarzschild value.
It is important to emphasize that the scale-dependent effects of the CWG parameters $\gamma$ and $k$ on gravitational redshift become observationally significant only at intermediate and large distance scales. Specifically, the linear term $\gamma r$ starts to influence gravitational redshift at galactic scales ($r\gtrsim 10$ kpc), where it plays a role in the flat rotation curve problem, producing measurable deviations from Schwarzschild behavior. The quadratic cosmological term $kr^{2}$ becomes relevant only at larger scales, particularly for distances beyond $r\gtrsim 30$ Mpc, where it dominates and corresponds to the regime of Hubble flow, in which $k$ effectively acts as the squared Hubble constant [87, 34]. At smaller scales, such as within stellar systems and at the BH horizon ($r\lesssim 1$ kpc), both $\gamma r$ and $kr^{2}$ terms remain negligible when compared to the dominant Newtonian $\beta/r$ contribution. In this regime, deviations from GR predictions are typically below $1\%$. The characteristic transition scales, where each CWG term becomes significant, are given by $r_{\gamma}\sim(\beta/\gamma)^{1/2}\sim 10$ kpc for the linear term and $r_{k}\sim\gamma/k\sim 30$ Mpc for the quadratic term. These transition scales establish a natural hierarchy, with the linear term dominating at intermediate (galactic) scales and the quadratic term becoming significant at larger (cosmological) scales, separated by approximately two orders of magnitude. This hierarchical behavior ensures that the effects of $\gamma r$ and $kr^{2}$ are negligible at BH horizon scales, as confirmed by both observational constraints and solar system tests of CWG [85, 84].
While we recognize the importance of understanding the gravitational redshift at intermediate and large scales, this manuscript focuses primarily on the regime near the BH horizon. However, as demonstrated in Tables 4 and 5, even for the moderate parameter values considered here, the CWG corrections produce substantial enhancements in the gravitational redshift that could potentially be detected through precision spectroscopic observations of compact objects, offering new avenues for testing alternative gravity theories.
8 Conclusion
@@
In this paper, we explored the thermodynamic and quantum features of BHs arising from the MK solution within the framework of CWG. Unlike the conventional Einstein-Hilbert paradigm based on second-order field equations, CWG employs a fourth-order action with local conformal symmetry, leading to fundamentally modified gravitational dynamics and significantly richer vacuum structures [70, 18]. The inclusion of additional potential terms, specifically the linear $\gamma r$ and quadratic $kr^{2}$ contributions in the metric function given by Eq. (9), enabled an alternative theoretical description of gravitational phenomena spanning both galactic and cosmological scales without requiring the introduction of dark matter components.
@@
Our analysis commenced with the application of the Hamilton-Jacobi tunneling formalism to compute the Hawking temperature of CWGBHs, as detailed in Section 3. This investigation revealed explicit contributions from all conformal parameters $\beta$, $\gamma$, and $k$, culminating in the temperature expression presented in Eq. (25). The resulting thermal spectrum deviates markedly from that of standard Schwarzschild-AdS BHs, particularly due to the influence of the linear potential term $\gamma r$, demonstrating the remarkable sensitivity of BH radiation to conformal corrections [21, 92]. Table 1 provided quantitative evidence of these effects using self-consistent horizon radii determined from $B(r_{h})=0$, ensuring that all thermodynamic quantities are evaluated at physically meaningful horizon values where the MK parameters and $r_{h}$ are properly coupled.
We subsequently incorporated quantum gravitational effects through the GUP framework, deriving the QC Hawking temperature expression that accounts for minimal length scale effects. This analysis demonstrated that the presence of a fundamental length scale systematically suppresses the thermal spectrum, particularly in the near-Planckian regime where $r_{h}\to l_{p}$. The resulting temperature suppression reflects a gravitational redshift in the emitted radiation and suggests the potential emergence of remnant-like structures during the final stages of BH evaporation [93, 119]. The theoretical consistency between GUP and CWG was established through the spontaneous symmetry breaking mechanism: while CWG possesses exact conformal symmetry at UV scales (forbidding dimensionful parameters like $G$), this symmetry is spontaneously broken at IR scales through scalar field vacuum expectation values, generating an effective Newton constant $G_{\text{eff}}$ and enabling the Planck scale to emerge dynamically [85, 86, 52]. The GUP corrections thus represent effective low-energy descriptions valid in the symmetry-broken phase. The investigation of GUP-influenced phase transitions reveals quantum gravitational effects that fundamentally modify the thermodynamic landscape of CWGBHs through restructuring of critical point architectures and thermal response mechanisms [31, 108, 26]. Temperature suppression arising from minimal length effects creates distinct quantum phases characterized by enhanced heat capacity complexity as demonstrated in Eq. (47), where modifications to thermal equilibrium conditions depend on both classical gravitational parameters ($\beta$, $\gamma$, $k$) and the fundamental quantum correction factor $\lambda$ [33, 11, 56]. The computational analysis presented in Figures 9 and 10, computed with parameters $\beta=2$, $\gamma=0.05$, and $k=1$ (yielding a classical critical radius $r_{\text{crit}}\approx 1.34$), validates theoretical predictions of monotonic critical radius increases and thermal response modifications that become increasingly pronounced as horizon dimensions approach Planck-scale regimes [13, 91, 24]. These quantum-induced phase modifications reshape gravitational stability landscapes through discrete energy spectra and modified dispersion relations, establishing CWG as a natural theoretical framework for investigating quantum gravitational phenomena where classical field theory descriptions become inadequate for capturing the complete spectrum of thermodynamic behavior [89, 19, 74, 126, 68, 14].
@@
In Section 5, we employed an exponentially corrected entropy model, as expressed in Eq. (34), to compute the complete set of QC thermodynamic potentials for CWGBHs. The choice of exponential over logarithmic corrections was motivated by their non-perturbative origin in microstate counting [32] and their relevance in the near-Planckian regime where GUP effects operate. We also established that the Wald entropy for the MK solution coincides with the Bekenstein-Hawking area law $S_{0}=\pi r_{h}^{2}$ in the symmetry-broken phase [149, 76]. These calculations yielded corrected expressions for internal energy (Eq. (36)), Helmholtz free energy (Eq. (38)), pressure (Eq. (40)), enthalpy (Eq. (42)), and heat capacity (Eq. (45)). The corrected internal energy (Eq. (36)) recovers the dimensionally consistent Schwarzschild limit $E=\beta\ln r_{h}$, while the corrected Gibbs free energy exhibits sign changes at $r_{h}=e^{5/6}$, signaling thermodynamic phase boundaries. We emphasized that these thermodynamic quantities are presented as parametric functions of $(r_{h},\gamma)$, which is standard practice for revealing functional dependence on horizon geometry, while Table 2 provides explicit numerical values computed at self-consistent horizons. Figure 3 illustrates the energy landscape, revealing how scale-dependent corrections become increasingly important in the small horizon radius regime. Most significantly, Figure 8 demonstrated the existence of thermodynamic phase transitions through the heat capacity analysis, showing distinct regions of negative and positive heat capacity that correspond to thermodynamically unstable and stable phases, respectively. The heat capacity divergence corresponds to second-order phase transitions at $kr_{h}^{3}+3\beta^{2}\gamma-2\beta=0$, which disappears entirely for $\gamma>\gamma_{\rm crit}=2/(3\beta)$.
@@
The JT effect analysis presented in Section 6 provided a sensitive probe of isenthalpic processes in CWGBH thermodynamics within the extended phase space formalism. The behavior of the JT coefficient $\mu_{J}$ revealed distinct cooling and heating phases that depend sensitively on both the horizon radius and the CWG parameters. The corrected $\mu_{J}$ given by Eq. (52) possesses a constant denominator $3(3\beta^{2}\gamma-2\beta)$, leading to a qualitatively different contour morphology compared to the original expression, with the cooling regime extending over a broader area at $r_{h}>r_{\rm inv}$. Table 3 demonstrated that at self-consistent event horizons, the BH resides in the heating regime ($\mu_{J}<0$) for all examined parameter combinations, with the JT inversion occurring at larger radii given by $r_{\text{inv}}=[(2\beta-3\beta^{2}\gamma)/k]^{1/3}$. The corresponding inversion temperature $T_{\rm inv}$ decreases with increasing $\gamma$, narrowing the cooling window until it vanishes at $\gamma_{\rm crit}=2/(3\beta)$. Figure 12 captured the QC phase structure, showing how the scale-dependent parameter $\gamma$ controls the location of JT inversion points and introduces increasingly complex thermal response behavior. The additional contour plot (Figure 13) provides a visualization of the heating-cooling phase boundary across the $(r_{h},\gamma)$ parameter space. These results demonstrated that CWG naturally accommodates rich phase structures that are absent in conventional GR treatments.
@@
Our investigation of gravitational redshift within CWGBH geometry, detailed in Section 7, revealed that the frequency shift of electromagnetic radiation emitted near these objects encodes valuable information about the underlying conformal metric structure. We derived the correct redshift formula $1+z=\sqrt{B(r_{o})/B(r_{e})}$ for an observer at finite radius $r_{o}$ and emitter at $r_{e}$, noting that the commonly used asymptotically flat approximation is invalid for the MK metric due to the divergent $\gamma r$ and $kr^{2}$ terms at large radii. Tables 4 and 5 quantify the CWG corrections to gravitational redshift, revealing enhancements exceeding 200% compared to Schwarzschild values for representative parameter choices. The derived redshift expression showed complex radial dependence arising from the interplay of all CWG parameters, providing potentially observable signatures that could serve as discriminating tests between CWG and standard GR through precision spectroscopic observations.
Finally, we acknowledge that while analytical solutions provide valuable insights into physical systems, obtaining exact closed-form solutions for CWGBHs, especially with the added $\gamma$ and $k$ parameters, proves mathematically challenging. The fourth-order Bach field equations lead to quintic polynomial conditions for event horizons, which, by Abel’s theorem, cannot be fully solved in a general analytical form. This limitation is not unique to our study but is a broader issue in higher-derivative gravity theories. Nonetheless, while we can derive approximate analytical solutions in specific limits or under simplifying assumptions, numerical methods are crucial for fully capturing the physical behaviors of the system. Our numerical analysis allows us to accurately explore the parameter space, providing precise results that reveal essential physical trends, such as thermodynamic phase transitions, critical phenomena, and scale-dependent corrections. These features are universal across a wide range of physically relevant parameters. The combination of analytical techniques with targeted numerical exploration not only provides deeper insights into the underlying physics of quantum-corrected CWGBHs but also sets the methodological standard for investigating theories beyond Einstein’s relativity in realistic astrophysical settings.
@@
Looking toward future research directions, our theoretical framework suggests several promising directions for extended investigation. First, the development of more sophisticated numerical techniques for solving the complete nonlinear CWG field equations would enable precise predictions of observational signatures, including gravitational wave emission patterns, accretion disk spectra, and electromagnetic radiation characteristics that could be compared with data from current and future astronomical facilities. Second, the extension of our thermodynamic analysis to rotating CWGBHs would provide crucial insights into the effects of angular momentum on QC thermal properties, potentially revealing new classes of phase transitions and instabilities relevant to realistic astrophysical scenarios. Third, observational studies involving gravitational lensing phenomena, particularly strong lensing effects around supermassive BHs, could provide direct constraints on the CWG parameters $\beta$, $\gamma$, and $k$ through comparison with high-resolution imaging data from facilities such as the Event Horizon Telescope. Fourth, analyzing gravitational wave signatures from CWGBHs during mergers might reveal unique waveform patterns different from those predicted by GR, providing opportunities for testing alternative gravity theories with LIGO/Virgo and future space detectors. Lastly, exploring CWG’s cosmological implications, especially its potential to explain dark matter and cosmic acceleration without exotic matter, could extend our BH-focused study. Thus, our future research aims to advance the understanding of quantum gravity and to provide methods for examining spacetime geometry in extreme astrophysical environments.
Acknowledgments
We are grateful to the anonymous referees for their constructive comments and suggestions, which led to significant improvements in the manuscript. We also thank Ercan Kılıçarslan and Heeseung Zoe for their constructive feedback and technical discussions that contributed to the refinement of our computational framework. İ. S. appreciates the academic support from EMU, TÜBİTAK, ANKOS, and SCOAP3, as well as the networking opportunities through COST Actions CA22113, CA21106, CA21136, CA23130, and CA23115, which facilitated collaborative research initiatives and international scientific exchanges essential for advancing our understanding of quantum gravitational phenomena in alternative gravity theories.
Declaration of generative AI and AI-assisted technologies in the manuscript preparation process
During the preparation of this work the authors used AI and AI-assisted technologies in order to refine language, improve grammar, and code . [^8] In accordance with Elsevier’s policy on the use of generative AI and AI-assisted technologies in scientific writing: https://www.elsevier.com/about/policies-and-standards/generative-ai-policies-for-journals. After using this tool, the authors reviewed and edited the content as needed and take full responsibility for the content of the published article. All intellectual content, analysis, and conclusions are the authors’ own.